# A Gaussian integral formula for the Hermite polynomials: Combinatorics, Asymptotics and Applications

### v1

@@ 4.6 The Airy limit
5 Applications to Dyson Brownian motion and non-intersection Brownian motions
In this section we introduce Dyson Brownian motion, which can be considered as the eigenvalue process of the matrix analogue of Brownian motion (Dyson Brownian motion) [Dys62] or as Brownian motion in a Weyl chamber, that is the fundamental chamber of the Type A reflection group [JO06]. We introduce determinantal point processes (DPP) and show that Dyson Brownian motion is a DPP with kernel given by the Hermite kernel from Definition 4.2. We can then appeal to the results from Sections 4.5 and 4.6 to obtain convergence of the empirical distribution to the semicircle law and of the top particle to the Tracy-Widom distribution.
@@ 5.1 Two constructions of DBM
$V(x)=\sum_{\alpha\in\Phi^{+}}\alpha\cdot x=\prod_{i<j}\left(x_{i}-x_{j}\right),$
where $\Phi^{+}=\left\{e_{i}-e_{j}\colon i<j\right\}$ is the positive root system of the reflection group $W(\Sigma)$. Note that the function $V$ is often called the Vandermonde determinant. By the general theory of $h$-transforms (see for example [Bia94]), the NIBM process is the diffusion
@@
It was shown by Cépa–Lépingle [CL97] that the SDE (66) has a unique solution and that $X_{1}(t)<X_{2}(t)<\ldots<X_{n}(t)$ almost surely for all $t>0$. On the other hand, the Hermitian Brownian motion is obtained by choosing $2n^{2}$ independent Brownian motions $B_{j\ell},\beta_{j\ell}$ ($1\leq j,\ell\leq n$) and then defining the $\mathbb{C}^{n\times n}$-valued process $H=(H_{j\ell})_{j,\ell=1}^{n}$ by
@@
As the name suggests, the matrix $H(t)$ is Hermitian for any $t>0$ and so its eigenvalues $X_{1}(t),...,X_{n}(t)$ are real. If we re-order them to be increasing, then [Dys62] they satisfy (66) with $\lambda_{j}(0)=0$.
@@
Figure 9: A Brownian watermelon
The proof of the following proposition is a straightforward calculation using the Karlin–McGregor formula. See Appendix B.1 for details.
@@
In order to prove Proposition 5.10, we appeal to Theorem 11.4.1 in [Bor11], which says informally that bi-orthogonal ensembles are determinantal point processes.
@@
Theorem 5.13 (Theorem 11.4.1 in [Bor11]).
@@ 5.3 Fluctuations of the top particle
Having seen that the empirical measure of the DBM process run up to time $\frac{1}{n}$ converges to the semicircle law, whose support is [2,2][-2,2], one can ask whether the top particle converges to the value $2$. We now show that it does and compute the fluctuations around this value in the process. See [AGZ09] and references therein. We will see that the limiting fluctuations are of order $n^{2/3}$ and are distributed according to the Tracy–Widom GUE distribution, which is closely related to the Airy function that already appeared in 4.6. For our calculations, it will be helpful to appeal to theory of Fredholm determinants. We start by giving a condensed introduction to this theory.
@@
One advantage of DPPs is that we can write observables in terms of determinants or Fredholm determinants. For background see for example [McK11]. A kernel on a a subset $D\subseteq\mathbb{R}^{d}$ is a map $K\colon D\times D\longrightarrow\mathbb{R}$. Under suitable integrability conditions, every kernel induces a linear map $T\colon L^{2}(D)\longrightarrow L^{2}(D)$ defined by $T(f)(x)=\int_{D}K(x,y)\,f(y)\,dy$. Note that every (suitable) kernel determines a unique operator,but that an operator may have more than one kernel that defines it. For a kernel $K$, we will write $T_{K}$ for the operator that it gives rise to.
@@
Note that $P_{(a,\infty)}KP_{(a,\infty)}$ denotes the operator with kernel $(x,y)\longmapsto 1_{(a,\infty)}(x)K(x,y)1_{(a,\infty)}(y)$. Proposition 5.17 follows from setting $\eta=-\mathbf{1}_{[a,\infty)}$ in the following, more general result that can be found in [Bor11], see (11.2.4). Note that if $S$ has a fixed number of points then it always has a top particle.
@@
We prove this for $\sigma=1$, the general result follows from Brownian scaling (Remark 5.2). Recall that the DBM process $X$ is determinantal with kernel $K^{\text{DBM}}_{n,\cdot}$ So, by Proposition 5.17 and a change of variables, the left hand side can be written as the Fredholm determinant $\det\left(1-\chi_{s}\widetilde{K}^{\text{DBM}}_{n}\chi_{s}\right)$ where
$\widetilde{K}^{\text{DBM}}_{n}(u,v)$ $=n^{-1/6}K_{n}\left(2\sqrt{n}+n^{-1/6}u,2\sqrt{n}+n^{-1/6}v\right),$ (76)

### v2

@@ 4.6 The Airy limit
5 Random matrix applications part I: Dyson Brownian motion
In this section we introduce Dyson Brownian motion, which can be considered as the eigenvalue process of the matrix analogue of Brownian motion (Dyson Brownian motion) [Dyson] or as Brownian motion in a Weyl chamber, that is the fundamental chamber of the Type A reflection group [JonesOConnell]. We introduce determinantal point processes (DPP) and show that Dyson Brownian motion is a DPP with kernel given by the Hermite kernel from Definition 4.2. We can then appeal to the results from Sections 4.5 and 4.6 to obtain convergence of the empirical distribution to the semicircle law and of the top particle to the Tracy-Widom distribution.
@@ 5.1 Two constructions of DBM
$V(x)=\sum_{\alpha\in\Psi^{+}}\alpha\cdot x=\prod_{i<j}\left(x_{i}-x_{j}\right),$
where $\Phi^{+}=\left\{e_{i}-e_{j}\colon i<j\right\}$ is the positive root system of the reflection group $W(\Sigma)$. Note that the function $V$ is often called the Vandermonde determinant. By the general theory of $h$-transforms (see for example [Biane1994]), the NIBM process is the diffusion
@@
It was shown by Cépa–Lépingle [CepaLepingle] that the SDE (73) has a unique solution and that $X_{1}(t)<X_{2}(t)<\ldots<X_{n}(t)$ almost surely for all $t>0$. On the other hand, the Hermitian Brownian motion is obtained by choosing $2n^{2}$ independent Brownian motions $B_{j\ell},\beta_{j\ell}$ ($1\leq j,\ell\leq n$) and then defining the $\mathbb{C}^{n\times n}$-valued process $H=(H_{j\ell})_{j,\ell=1}^{n}$ by
@@
As the name suggests, the matrix $H(t)$ is Hermitian for any $t>0$ and so its eigenvalues $X_{1}(t),...,X_{n}(t)$ are real. If we re-order them to be increasing, then [Dyson] they satisfy (73) with $\lambda_{j}(0)=0$.
@@
Figure 10: A Brownian watermelon with $n=6$ non-intersecting Brownian bridges.
The proof of the following proposition is a straightforward calculation using the Karlin–McGregor formula. See Appendix C.1 for details.
@@
$\frac{1}{\sqrt{t}}\sum_{j=0}^{n}\Psi_{j}(x/\sqrt{t})\Psi_{j}(y\sqrt{t})$
In order to prove Proposition 5.10, we appeal to Theorem 11.4.1 in [Borodin-HbRMT], which says informally that bi-orthogonal ensembles are determinantal point processes.
@@
Theorem 5.13 (Theorem 11.4.1 in [Borodin-HbRMT]).
@@ 5.3 Fluctuations of the top particle
Having seen that the empirical measure of the DBM process run up to time $\frac{1}{n}$ converges to the semicircle law, whose support is [2,2][-2,2], one can ask whether the top particle converges to the value $2$. We now show that it does and compute the fluctuations around this value in the process. Figure 11 illustrates this. See also [AGZ] and references therein. We will see that the limiting fluctuations are of order $n^{2/3}$ and are distributed according to the Tracy–Widom GUE distribution, which is closely related to the Airy function that already appeared in 4.6.
Figure 11: A sample of the Dyson Brownian motion process when $n=36$. The trajectory of the top particle, $X_{1}(\sigma/n)$ is bolded. The LLN behaviour at $y=2\sqrt{\sigma}$ is illustrated, and a 90% confidence interval using the result of Proposition 5.19, along with the quantiles for the Tracy-Widom distribution $F_{2}(-3.19)=0.05$, $F_{2}(-0.23)=0.95$ computed in [Chiani].
For our calculations, it will be helpful to appeal to theory of Fredholm determinants. We start by giving a condensed introduction to this theory.
@@
One advantage of DPPs is that we can write observables in terms of determinants or Fredholm determinants. For background see for example [McKean]. A kernel on a a subset $D\subseteq\mathbb{R}^{d}$ is a map $K\colon D\times D\longrightarrow\mathbb{R}$. Under suitable integrability conditions, every kernel induces a linear map $T\colon L^{2}(D)\longrightarrow L^{2}(D)$ defined by $T(f)(x)=\int_{D}K(x,y)\,f(y)\,dy$. Note that every (suitable) kernel determines a unique operator,but that an operator may have more than one kernel that defines it. For a kernel $K$, we will write $T_{K}$ for the operator that it gives rise to.
@@
Note that $P_{(a,\infty)}KP_{(a,\infty)}$ denotes the operator with kernel $(x,y)\longmapsto 1_{(a,\infty)}(x)K(x,y)1_{(a,\infty)}(y)$. Proposition 5.17 follows from setting $\eta=-\mathbf{1}_{[a,\infty)}$ in the following, more general result that can be found in [Borodin-HbRMT], see (11.2.4). Note that if $S$ has a fixed number of points then it always has a top particle.
@@
Recall that the DBM process $X$ is determinantal with kernel $K^{\text{DBM}}_{n,\cdot}$. So, by Proposition 5.17 and a change of variables, the left hand side can be written as the Fredholm determinant $\det\left(1-\chi_{s}\widetilde{K}^{\text{DBM}}\chi_{s}\right)$ where
$\widetilde{K}^{\text{DBM}}_{n}(u,v)$ $=n^{-2/3}\sigma^{-1/2}K^{\text{DBM}}_{n,\sigma/n}\left(2\sqrt{\sigma}+\frac{u}{n^{2/3}\sigma^{1/2}},2\sqrt{\sigma}+\frac{v}{n^{2/3}\sigma^{1/2}}\right)$ (83)
$=n^{-2/3}\sigma^{-1/2}\frac{\sqrt{n}}{\sigma}K_{n}\left(2\sqrt{n}+n^{-1/6}\frac{u}{\sigma},2\sqrt{n}+n^{-1/6}\frac{v}{\sigma}\right)$ (84)
$=n^{-1/6}\frac{1}{\sigma}K_{n}\left(2\sqrt{n}+n^{-1/6}\frac{u}{\sigma},2\sqrt{n}+n^{-1/6}\frac{v}{\sigma}\right)$ (85)